OpenStudy (anonymous):

What is the solution set for |3x + 12|+ 5 = 23?

OpenStudy (anonymous):

x=2,-10

OpenStudy (anonymous):

ahh i just got that :) 3*x + 12 + 5 = 23 3*x = 23 - 17 = 6 x = 2 - (3*x + 12) + 5 = 23 - 3*x - 12 + 5 = 23 - 7 - 23 = 3*x 3*x = - 30 x = - 10